OpenStudy (anonymous):

How can x + 4 = 3x + 12 be set up as a system of equations?

OpenStudy (anonymous):

i need help

OpenStudy (anonymous):

@ImNotGoodAtMath

OpenStudy (anonymous):

@satellite73

OpenStudy (tylerd):

solve for x first im guessing

jimthompson5910 (jim_thompson5910):

For the given equation, the two sides are equal to one another. So that means they are equal to the same variable, say y. So y = x+4 y = 3x + 12 notice if we were to start with y = 3x + 12 and replace y with x+4 (valid because y = x+4), we'd get the original equation back again.

OpenStudy (anonymous):

thamks